Ideal waist size. According to Dr. Aaron R. Folsom of the University Of Minnesota School Of Public Health, your maximum ideal waist size is directly proportional to your hip size. For a woman with 40-inch hips, the maximum ideal waist size is 32 inches. What is the maximum ideal waist size for a woman with 35-inch hips?
Answer by stanbon(75887)

maximum ideal waist size is directly proportional to your hip size. For a woman with 40-inch hips, the maximum ideal waist size is 32 inches. What is the maximum ideal waist size for a woman with 35-inch hips?
----
Equation:
w = kh
---
Solve for "k" using "For a woman with 40-inch hips, the maximum ideal waist size is 32 inches".
---
32 = k40
k = 4/5
----
Equation for your problem:
w = (4/5)h
---
What is the maximum ideal waist size for a woman with 35-inch hips?
---
w = (4/5)35
w = 4*7
w = 28 inches
========================
Cheers,
Stan H.